# TVR-Ranking: A Dataset for Ranked Video Moment Retrieval with Imprecise Queries

### v1

@@ 6 Conclusion and Limitations
Supplementary Material for TVR-Ranking:
A Dataset for Ranked Video Moment Retrieval with Imprecise Queries
@@ Appendix A Character Name Substitution
Furthermore, we add a verification process to avoid potential errors in ChatGPT’s responses and to ensure output quality. The verification contains two steps. First, we assess the semantic alignment of the query before and after word substitution, using SimCSE Gao et al. (2021) to calculate the sentence similarity. Second, we use Flair Akbik et al. (2019) to check if named entity recognition (NER) can be detected from the replaced version. Outputs deemed of inadequate quality are reprocessed by ChatGPT with a different random seed and subsequently reviewed by annotators.
@@ B.1 Annotation Guideline
The moment precisely matches the search query, though there are a few noticeable mismatches. The action within the moment only conveys a part of the query information. There is a large gap in visual and temporal completeness.
@@ B.2 Annotation Setup and Annotators
We received the number of 227,808 raw annotations. After cleaning and merging, we obtained the number of 94,442 annotated moments for 3,281 queries with annotation consensus.
@@ B.5 Annotation Analysis: Case Study of Example Queries
We clarify two points here. First, each query is a moment caption originated from the TVR dataset, hence there exists at least one candidate moment whose caption is identical to the query, i.e., $sim(q,m.c)=1.0$. Second, the moment captions in Figure 6 are provided for reference purposes here. The captions of candidate moments are not shown to the annotators during the annotation process. Annotators judge the level of relevance purely based on moment’s visual content to the query. Further, moment captions in the dataset shall only be used as queries, and not as additional information available in source videos. In a typical RMVR task, the videos are not segmented into moments with high-quality captions.
@@ Appendix C Experiments and Case Study on Evaluation
XML Lei et al. (2020) 1 0.1017 0.0999 0.0767 0.0729 0.0283 0.0312
20 0.1539 0.1159 0.0441 0.1547 0.1170 0.0464
40 0.1635 0.1215 0.0472 0.1643 0.1242 0.0504
CONQUER Hou et al. (2021) 1 0.1167 0.1035 0.0981 0.0814 0.0631 0.0543
20 0.2193 0.2119 0.1981 0.1954 0.1489 0.1434
40 0.2304 0.2071 0.2083 0.1909 0.1556 0.1382
ReLoCLNet Zhang et al. (2021) 1 0.1299 0.1305 0.1177 0.1169 0.0756 0.0814
20 0.2739 0.2626 0.2522 0.2442 0.1798 0.1786
40 0.2808 0.2785 0.2636 0.2614 0.1885 0.1940
$NDGC@20$
@@
XML Lei et al. (2020) 1 0.1015 0.0998 0.0751 0.0731 0.0289 0.0313
20 0.1670 0.1287 0.0530 0.1554 0.1155 0.0478
40 0.1767 0.1356 0.0582 0.1624 0.1228 0.0514
CONQUER Hou et al. (2021) 1 0.1143 0.1029 0.0956 0.0811 0.0633 0.0539
20 0.2014 0.1946 0.1817 0.1777 0.1363 0.1313
40 0.2068 0.1919 0.1875 0.1768 0.1428 0.1291
ReLoCLNet Zhang et al. (2021) 1 0.1234 0.1236 0.1110 0.1100 0.0712 0.0770
20 0.2615 0.2545 0.2424 0.2373 0.1752 0.1762
40 0.2715 0.2700 0.2552 0.2541 0.1850 0.1913
$NDGC@40$
@@
XML Lei et al. (2020) 1 0.1070 0.1064 0.0796 0.0788 0.0319 0.0337
20 0.1566 0.1184 0.0505 0.1687 0.1292 0.0567
40 0.1636 0.1250 0.0539 0.1761 0.1364 0.0607
CONQUER Hou et al. (2021) 1 0.1186 0.1082 0.0984 0.0853 0.0657 0.0569
20 0.2006 0.1941 0.1798 0.1753 0.1365 0.1299
40 0.2052 0.1892 0.1832 0.1717 0.1392 0.1250
ReLoCLNet Zhang et al. (2021) 1 0.1273 0.1274 0.1138 0.1133 0.0730 0.0793
20 0.2701 0.2628 0.2509 0.2455 0.1847 0.1848
40 0.2787 0.2766 0.2625 0.2613 0.1935 0.1988
In our experiments, we adapt three VCMR models to RVMR and evaluate them on the TRV-Ranking: XML Lei et al. (2020), CONQUER Hou et al. (2021), and ReLoCLNet Zhang et al. (2021).
The main change in the adaptations is the introduction of weight to the training loss based on query-moment similarity, i.e., $sim(q,m.c)$, recognizing that moments vary in relevance to a query. There is no such loss in the VCMR task setting because there is exactly one ground truth in VCMR. This weight aims to diminish the influence of less relevant moments on model training by adjusting the loss. In terms of features, we follow the original implementation and use both subtitle and video features extracted from the source video for all three models. The query features in our implementation are extracted using BERT Devlin et al. (2018).
Note that, because VCMR was directly extended from VMR, there is a train/test split in the source videos. That is, there is a set of queries and source videos for training, and another set for testing. However, the VCMR task is similar to web search, except the result documents are video moments. In the VCMR task setting, both training and test queries search for moments from a common large pool of source videos. As long as there are no duplicates between training and test queries, there is no data leakage issue.
@@ C.1 Baselines and Implementation Details
Cross-modal Moment Localization (XML) Lei et al. (2020). The XML model was proposed alongside the TVR dataset. The motivation behind XML is to consider both video and subtitle information when retrieving moments, as some queries in TVR are based on subtitles. The model integrates video and subtitle features as context information and conducts retrieval on this context to achieve more accurate recall. During the retrieval stage, the authors use matrix multiplication to compute the confidence score, enhancing retrieval efficiency. Additionally, they designed another branch specifically to predict the start and end times of moments. The XML model is licensed under the MIT license.
The video features in XML combine visual features from ResNet He et al. (2016) and temporal features from I3D Carreira and Zisserman (2017). The subtitle features are extracted via RoBERTa Liu et al. (2019). For the loss, we consider two factors. First, we multiply the similarity $\text{sim}(q,m.c)$ (ranging from 0.0 to 1.0) in the original loss as a decay coefficient. We expect moments that are more likely to be relevant to have a greater impact on the model during training. Additionally, XML employs a video-level contrastive loss, pulling positive videos closer and pushing negative videos farther away. However, in our task setting, there could be multiple positive videos within a mini-batch for one query, because in the pseudo training set all top $N$ videos can be regarded as positive samples. To address the issue, we introduce a positive pair mask Khosla et al. (2020). This mask assigns a value of one to positions where the query and video are related in the pseudo training set, and zeros out other positions. When computing the loss, we sum the scores masked by the positive pair mask instead of using only the diagonal elements as the nominator.
Retrieval and Localization Network with Contrastive Learning (ReLoCLNet) Zhang et al. (2021). The structure of ReLoCLNet is similar to XML. Both models integrate video and subtitle features as context, separately encode query and context embeddings, and use video-level contrastive loss. The key difference is that ReLoCLNet adds a frame-level contrastive loss to highlight the ground truth region and improve recall accuracy. The model is under the MIT license.
@@
CONtextual QUery-awarE Ranking (CONQUER) Hou et al. (2021). CONQUER uses a different framework compared to the other two models. It first retrieves candidate videos from the corpus and then conducts moment localization within those candidate videos.
CONQUER uses video features that concatenate the SlowFast Feichtenhofer et al. (2019) and ResNet models, while subtitle features are extracted using RoBERTa Liu et al. (2019). Additionally, CONQUER leverages video retrieval results from HERO Li et al. (2020). We continue to use these results when training on our TVR-Ranking dataset. Note that, because the HERO results are obtained from the TVR dataset, there could be a data leak issue in our task setting. However, this issue is negligible for two reasons: (i) the queries used in our setting is imprecise query with query re-written, and (ii) a query has multiple ground truth moments in our task setting, which was not annotated in the original TVR dataset.
For all three models, we set the batch size to 128 and the learning rate to 0.0001 with a warmup phase. For the $N=1$ pseudo training set, we trained the models for 2000 epochs, evaluating the validation set every 10 epochs. For the $N=20$ and $N=40$ training sets, we trained for 50 and 25 epochs, respectively, and evaluated the validation set 2 and 4 times per epoch, respectively. We adhered to the original model parameters and did not use non-maximum suppression for post-processing the outputs. All models were adapted to the same PyTorch version (2.2.1) with CUDA 12.1. The experiments were conducted on a single NVIDIA V100 32GB GPU.
@@ C.2 Experiment Results
Table 6 reports the full experiment results. The three sets of experiments are conducted with different $N=\{1,20,40\}$ values, and evaluated by $NDCG@K,IoU\geq\mu$, for $K=\{10,20,40\}$ and $\mu=\{0.3,0.5,0.7\}$. The parameter search is based on the best results on the validation set with $NDCG@20,IoU\geq 0.5$. We also ran the ReLoCLNet model three times with different seeds to report the error bars, shown in 7.
Table 7: Performance of the ReLoCLNet under different seeds.
$IoU\geq 0.3$ $IoU\geq 0.5$ $IoU\geq 0.7$
seed val test val test val test
$NDGC@10$
$s_{1}$ 0.2739 0.2626 0.2522 0.2442 0.1798 0.1786
$s_{2}$ 0.2772 0.2676 0.2495 0.2439 0.1714 0.1738
$s_{3}$ 0.2754 0.2778 0.2564 0.2589 0.1829 0.1910
$NDGC@20$
$s_{1}$ 0.2615 0.2545 0.2424 0.2373 0.1752 0.1762
$s_{2}$ 0.2698 0.2606 0.2441 0.2387 0.1715 0.1732
$s_{3}$ 0.2668 0.2700 0.2488 0.2525 0.1819 0.1887
$NDGC@40$
$s_{1}$ 0.2808 0.2785 0.2636 0.2614 0.1885 0.1940
$s_{2}$ 0.2819 0.2716 0.2554 0.2490 0.1841 0.1831
$s_{3}$ 0.2741 0.2771 0.2564 0.2599 0.1913 0.1971
To explore the difference between the VCMR task and our RVMR task, we compare these models’ performance on two tasks. The results on VCMR are summarized in Table 8, taken from the original papers  Lei et al. (2020) Hou et al. (2021) Zhang et al. (2021). For the VCMR task, CONQUER demonstrates the best performance on the VCMR task, surpassing ReLoCLNet by 3.61% with $IoU\geq 0.7$ and $R=1$, and by 8.34% with $IoU\geq 0.7$ and $R=11$. However, CONQUER falls behind ReLoCLNet on the RVMR task, with $NDCG@20$ lower by 0.0781, 0.0773, and 0.0622 for $IoU\geq\{0.3,0.5,0.7\}$, respectively. XML performed well on the VCMR task as well, achieving 44.44% with $IoU\geq 0.5$ and $R=100$, slightly outperforming ReLoCLNet’s 44.10%. However, it showed much poorer performance on the RVMR task. The discrepancy implies that the abilities required for the RVMR task differ from those for the VCMR task. Although VCMR models can be easily adapted, directly applying them to an RVMR application may not be appropriate. Thus, designing a new model tailored specifically to our RVMR task is necessary.
@@ C.3 Case study of Predictions for Sample Queries
To provide a more intuitive understanding of our TVR-Ranking task, we present three cases with different scores. These queries from the test set and the results from ReLoCLNet are listed in Figure 7. The predicted results are measured by $NDCG@10,IoU\geq 0.3$. Ground truth moments and predicted moments (with relevance levels indicated on the side), are presented in the left and right columns respectively. The grey videos in the ground truth indicate moments that were not found. The grey checkered videos in the results indicate incorrect predictions.
The first case is a low-score example, where we can see that only two moments were correctly matched within the predictions. The second case has an average score, with four relevant moments found, and the model correctly ranked a highly relevant moment at the top. The third case achieved a high score, with the top four predictions being correct, demonstrating excellent performance for our task.
(a) Result = 0.0783.
(b) Result= 0.2712.
(c) Result=0.6572.
Figure 7: Three example queries, along with their ground truth moments and predicted moments (with relevance levels indicated on the side), are presented. The predicted results are measured by $NDCG@10,IoU\geq 0.3$. The grey videos in the ground truth indicate moments that were not found, and the grey checkered videos in the results indicate incorrect predictions.

### v2

@@ Appendix A Character Name Substitution
Furthermore, we add a verification process to avoid potential errors in ChatGPT’s responses and to ensure output quality. The verification contains two steps. First, we assess the semantic alignment of the query before and after word substitution, using SimCSE [18] to calculate the sentence similarity. Second, we use Flair [19] to check if named entity recognition (NER) can be detected from the replaced version. Outputs deemed of inadequate quality are reprocessed by ChatGPT with a different random seed and subsequently reviewed by annotators.
@@ B.1 Annotation Guideline
The moment matches the search query, though there are a few noticeable mismatches. The action within the moment only conveys a part of the query information. There is a large gap in visual and temporal completeness.
@@ B.2 Annotation Setup and Annotators
We received the number of 227,808 raw annotations from all annotators. After cleaning and merging, we obtained the number of 94,442 annotated moments for 3,281 queries with annotation consensus.
@@ B.5 Annotation Analysis: Case Study of Example Queries
We clarify two points here. First, each query is a moment caption originated from the TVR dataset, hence there exists at least one candidate moment whose caption is identical to the query, i.e., $sim(q,m.c)=1.0$. Second, the moment captions in Figure 6 are provided for reference purposes here. The captions of candidate moments are not shown to the annotators during the annotation process. Annotators judge the level of relevance purely based on moment’s visual content to the query. Further, moment captions in the dataset shall only be used as queries, and not as additional information available in source videos. In a RMVR task, the videos are not segmented into moments and such high-quality captions do not exist as well.
@@ Appendix C Experiments and Case Study on Evaluation
XML [6] 1 0.1016 0.0917 0.0747 0.0660 0.0244 0.0268
20 0.2226 0.2135 0.1623 0.1567 0.0580 0.0627
40 0.2002 0.2044 0.1461 0.1502 0.0541 0.0589
CONQUER [21] 1 0.0999 0.0859 0.0844 0.0709 0.0530 0.0512
20 0.2406 0.2249 0.2222 0.2104 0.1672 0.1517
40 0.2450 0.2219 0.2262 0.2085 0.1670 0.1515
ReLoCLNet [22] 1 0.1575 0.1525 0.1358 0.1349 0.0908 0.0916
20 0.3751 0.3751 0.3407 0.3397 0.2316 0.2338
40 0.4339 0.4353 0.3984 0.3986 0.2693 0.2807
$NDCG@20$
@@
XML 1 0.1010 0.0923 0.0737 0.0662 0.0258 0.0269
20 0.2331 0.2243 0.1700 0.1650 0.0627 0.0664
40 0.2114 0.2167 0.1530 0.1590 0.0583 0.0635
CONQUER 1 0.0952 0.0835 0.0808 0.0687 0.0526 0.0484
20 0.2130 0.1995 0.1976 0.1867 0.1527 0.1368
40 0.2183 0.1968 0.2022 0.1851 0.1524 0.1365
ReLoCLNet 1 0.1504 0.1439 0.1303 0.1269 0.0866 0.0849
20 0.3815 0.3792 0.3462 0.3427 0.2381 0.2386
40 0.4418 0.4439 0.4060 0.4059 0.2787 0.2877
$NDCG@40$
@@
XML 1 0.1077 0.1016 0.0775 0.0727 0.0273 0.0294
20 0.2580 0.2512 0.1874 0.1853 0.0705 0.0753
40 0.2408 0.2432 0.1740 0.1791 0.0666 0.0720
CONQUER 1 0.0974 0.0866 0.0832 0.0718 0.0557 0.0510
20 0.2029 0.1906 0.1891 0.1788 0.1476 0.1326
40 0.2080 0.1885 0.1934 0.1775 0.1473 0.1323
ReLoCLNet 1 0.1533 0.1489 0.1321 0.1304 0.0878 0.0869
20 0.4039 0.4031 0.3656 0.3648 0.2542 0.2567
40 0.4725 0.4735 0.4337 0.4337 0.3015 0.3079
In our experiments, we adapt three VCMR models to RVMR and evaluate them on the TRV-Ranking: XML [6], CONQUER [21], and ReLoCLNet [22]. The main change in the adaptations is the introduction of weight to the training loss based on query-moment similarity, i.e., $sim(q,m.c)$, recognizing that moments vary in relevance to a query. There is no such loss in the VCMR task setting because there is exactly one ground truth in VCMR. This weight aims to diminish the influence of less relevant moments on model training by adjusting the loss. In terms of features, we follow the original implementation and use both subtitle and video features extracted from the source video for all three models. The query features in our implementation are extracted using BERT [23].
Note that, because VCMR was directly extended from VMR, there is a train/test split in the source videos. That is, there is a set of queries and source videos for training, and another set for testing in VCMR. However, the RVMR task is similar to web search, except the result documents are video moments. In the RVMR task setting, both training and test queries search for moments from a common large pool of source videos. As long as there are no duplicates between training and test queries, there is no data leakage issue.
@@ C.1 Baselines and Implementation Details
Cross-modal Moment Localization (XML) [6]. The XML model was proposed alongside the TVR dataset. The motivation behind XML is to consider both video and subtitle information when retrieving moments, as some queries in TVR are based on subtitles. The model integrates video and subtitle features as context information and conducts retrieval on this context to achieve more accurate recall. During the retrieval stage, the authors use matrix multiplication to compute the confidence score, enhancing retrieval efficiency. Additionally, they designed another branch specifically to predict the start and end times of moments. The XML model is licensed under the MIT license.
The video features in XML combine visual features from ResNet [24] and temporal features from I3D [25]. The subtitle features are extracted via RoBERTa [26]. For the loss, we consider two factors. First, we multiply the similarity $\text{sim}(q,m.c)$ (ranging from 0.0 to 1.0) in the original loss as a decay coefficient. We expect moments that are more likely to be relevant to have a greater impact on the model during training. Additionally, XML employs a video-level contrastive loss, pulling positive videos closer and pushing negative videos farther away. However, in our task setting, there could be multiple positive videos within a mini-batch for one query, because in the pseudo training set all top $N$ videos can be regarded as positive samples. To address the issue, we introduce a positive pair mask [27]. This mask assigns a value of one to positions where the query and video are related in the pseudo training set, and zeros out other positions. When computing the loss, we sum the scores masked by the positive pair mask instead of using only the diagonal elements as the nominator.
Retrieval and Localization Network with Contrastive Learning (ReLoCLNet) [22]. The structure of ReLoCLNet is similar to XML. Both models integrate video and subtitle features as context, separately encode query and context embeddings, and use video-level contrastive loss. The key difference is that ReLoCLNet adds a frame-level contrastive loss to highlight the ground truth region and improve recall accuracy. The model is under the MIT license.
@@
CONtextual QUery-awarE Ranking (CONQUER) [21]. CONQUER uses a different framework compared to the other two models. It first retrieves candidate videos from the corpus and then conducts moment localization within those candidate videos.
CONQUER uses video features that concatenate the SlowFast [28] and ResNet models, while subtitle features are extracted using RoBERTa [26]. Additionally, CONQUER leverages video retrieval results from HERO [29]. We continue to use these results when training on our TVR-Ranking dataset. Note that, because the HERO results are obtained from the TVR dataset, there could be a data leak issue in our task setting. However, this issue is negligible for two reasons: (i) the queries used in our setting is imprecise query with query re-written, and (ii) a query has multiple ground truth moments in our task setting, which was not annotated in the original TVR dataset.
For all three models, we set the learning rate to 0.0001 with a warmup phase. For the $N=1$ pseudo training set, we trained the models for 4000 epochs, evaluating the validation set every 20 epochs. For the $N=20$ and $N=40$ training sets, we trained for 200 and 100 epochs, respectively, and evaluated the validation set 1 and 2 times per epoch, respectively. We used early stopping, set to trigger if performance did not improve after 10 evaluations. We adhered to the original model parameters and did not use non-maximum suppression for post-processing the outputs. All models were adapted to the same PyTorch version (2.2.1) with CUDA 12.1. The experiments were conducted on a single NVIDIA V100 32GB GPU.
@@ C.2 Experiment Results
Table 6 reports the full results, conducted with different $N=\{1,20,40\}$ values, and evaluated by $NDCG@K,IoU\geq\mu$, for $K=\{10,20,40\}$ and $\mu=\{0.3,0.5,0.7\}$. The parameter search is based on the best results on the validation set with $NDCG@20,IoU\geq 0.5$. We have also conducted three runs of the ReLoCLNet model using different seeds with the top 40 pseudo training set. As reported in Table 7, the results are highly consistent with very small standard deviations
Table 7: Average $NDCG@K,IoU=\mu$ scores and their standard deviations of the ReLoCLNet model. Results are obtained from three runs of the model using different seeds, with the top 40 pseudo training set.
$NDCG$ $IoU\geq 0.3$ $IoU\geq 0.5$ $IoU\geq 0.7$
@ val test val test val test
10 $0.4287\pm 0.0048$ $0.4347\pm 0.0008$ $0.3943\pm 0.0038$ $0.3982\pm 0.0009$ $0.2696\pm 0.0029$ $0.2809\pm 0.0006$
20 $0.4359\pm 0.0053$ $0.4425\pm 0.0010$ $0.4010\pm 0.0045$ $0.4050\pm 0.0008$ $0.2761\pm 0.0033$ $0.2871\pm 0.0010$
40 $0.4668\pm 0.0050$ $0.4724\pm 0.0008$ $0.4290\pm 0.0041$ $0.4326\pm 0.0008$ $0.2983\pm 0.0041$ $0.3077\pm 0.0007$
To explore the difference between the VCMR task and our RVMR task, we compare these models’ performance on both tasks. The results on VCMR are summarized in Table 8, taken from the original papers  [6, 21, 22]. For the VCMR task, CONQUER demonstrates the best performance on the VCMR task, surpassing ReLoCLNet by 3.61% with $IoU\geq 0.7$ and $R=1$, and by 8.40% with $IoU\geq 0.7$ and $R=10$. However, CONQUER falls behind ReLoCLNet on the RVMR task for test set, with $NDCG@20$ lower by 0.2471, 0.2208, 0.1512 for $IoU\geq{0.3,0.5,0.7}$ respectively when using top 40 pseudo training set. XML performs well on the VCMR task as well, achieving 44.44% with $IoU\geq 0.5$ and $R=100$, slightly outperforming ReLoCLNet’s 44.10%. However, it showed much poorer performance on the RVMR task. The discrepancy implies that the abilities required for the RVMR task differ from those for the VCMR task. Although VCMR models can be easily adapted, directly applying them to an RVMR application may not be appropriate. Thus, designing a new model tailored specifically to our RVMR task is necessary.
@@ C.3 Case study of Predictions for Sample Queries
To provide a more intuitive understanding of the RVMR task and evaluation, we present three samples with different scores in Figure 7. These queries are from the test set and the results are by ReLoCLNet when using top 40 pseudo training set. The predicted results are measured by $NDCG@10,IoU\geq 0.3$. Ground truth moments and the predicted moments (with relevance levels indicated on the side) are presented in the left and right columns respectively, for each query. To clearly show the mapping of ground truth and the predicted moments, we mask the ground truth moments that are missed by the model. The masked videos along the prediction column means incorrect predictions i.e., not matching any ground truth.
The first case shows a high $NDCG$ score as the model successfully finds the most relevant moment for the query, along with some other relevant moments. The second case shows a similar sets of predicted relevant moments. However, this query has a few highly relevant ground truth moments, leading to a $NDCG$ score close to 0.5. The third case is low-score example where the model retrieves three relatively low relevant moments, missing the highly relevant ones.
(a) Result = 0.6167.
(b) Result= 0.4675.
(c) Result=0.1523.
Figure 7: Three example queries, along with their ground truth moments (on the left) and the predicted moments (on the right) with relevance levels indicated on their sides. The predicted results are measured by $NDCG@10,IoU\geq 0.3$. To clearly show the mapping of ground truth and the predicted moments, we mask the ground truth that are missed by the model. The masked videos along the prediction column means incorrect predictions i.e., not matching any ground truth.